OpenStudy (haleyelizabeth2017):

A 410g cylinder of brass is heated to 95.0°C and placed in a calorimeter containing 335g of water at 25.0°C. The water is stirred and its highest temperature is recorded at 32.0°C. From the thermal energy gained by the water, determine the specific heat of brass. The specific heat of water is 4.18 J/g · °C

OpenStudy (leonardo0430):

what would be the question?

OpenStudy (haleyelizabeth2017):

I have to find the specific heat of brass....I just need help getting started then someone to check my answer after I finish...

OpenStudy (matt101):

Whatever heat is lost by the brass is gained by the water. That means: \[m_{brass} c_{brass} \Delta T_{brass}=m_{water} c_{water} \Delta T_{water}\]

OpenStudy (haleyelizabeth2017):

oh

OpenStudy (haleyelizabeth2017):

mass is m and temperature? is t, correct?

OpenStudy (leonardo0430):

basically what matt said use the equation to determine the end product

OpenStudy (leonardo0430):

correct

OpenStudy (haleyelizabeth2017):

So I just plug everything and solve for c

OpenStudy (haleyelizabeth2017):

Okay....give me a second....would you mind checking my answer once I get it?

OpenStudy (leonardo0430):

yes so just plug in the equation, sure ill check it

OpenStudy (haleyelizabeth2017):

Okay, thank you

OpenStudy (haleyelizabeth2017):

Do i do the final temperature or beginning temperature or temp difference?

OpenStudy (leonardo0430):

the beginning

OpenStudy (haleyelizabeth2017):

Okay....

OpenStudy (matt101):

∆T is the temperature difference! Final temperature minus initial temperature

OpenStudy (haleyelizabeth2017):

Okay.

OpenStudy (haleyelizabeth2017):

One second while I make a few adjustments ;)

OpenStudy (haleyelizabeth2017):

Do I do the same to the brass because we know that it lost 7 degrees....since it transferred to the water

OpenStudy (leonardo0430):

yes do the same :)

OpenStudy (haleyelizabeth2017):

So what I did was: (410)(88)c=(335)(4.18)(7) 36,080c=9,802.1 c=0.27 Did I do it right?

OpenStudy (matt101):

NO @leonardo0430 ! Just because the water GAINED 7 degrees, doesn't mean the brass LOST 7 degrees. This would imply that specific heats of both substances are the same, which is not the case. Eventually the water and brass reach thermal equilibrium when they are BOTH the SAME TEMPERATURE. If the highest (i.e. final) temperature of the water is 32 degrees, then that is also the final temperature of the brass! If the brass were still 88 degrees when the water is 32 degrees, thermal equilibrium has not been reached. The brass would continue transferring heat to the water, and the temperature of the water would continue to rise.

OpenStudy (haleyelizabeth2017):

Ugh this is so confusing :(

OpenStudy (haleyelizabeth2017):

What do I do then @matt101?

OpenStudy (haleyelizabeth2017):

Ohhh! 410*63?

OpenStudy (haleyelizabeth2017):

Then I would get about 0.380....

OpenStudy (matt101):

The CHANGE in temperature is what we want. For water: final temp = 32, initial temp = 25, change in temp = 32-25 = 7 For brass: final temp = 32, initial temp = 95, change in temp = 32-95 = -63

OpenStudy (matt101):

Yes!

OpenStudy (haleyelizabeth2017):

okay.....

OpenStudy (haleyelizabeth2017):

Thank you

OpenStudy (matt101):

Btw, I should add to the equation I wrote out at the beginning that there should be a - on one side, because the heat gained by one side is lost by the other (on side is a positive change the other is a negative change). Having this negative in the equation will cancel the negative from the change in temperature so that your c will be positive. And yes, c=0.38!

OpenStudy (haleyelizabeth2017):

Okay..... :)